# Asymmetric Graph Error Control with Low Complexity in Causal Bandits

### v1

@@ III-B Edge-weighted Mutual Information
As a general measure of dependence, mutual information (MI), denoted by $I(\cdot)$, is widely utilized for causal inference (see e.g. [29, 30]). However, since the sub-graphs are treated separately, pure MI based sub-graph learning with limited data suffers significantly from stochastic disturbance. In this subsection, We explain this issue in detail and propose the edge-weighted mutual information measure as a solution.
@@
With the estimated edge-weight matrix denoted by $\hat{\bm{B}}_{\bm{a}}^{t}$, the estimated flow-weight matrix is defined as
$\hat{\bm{C}}_{\bm{a}}^{t}\doteq\bigl{(}\bm{I}-\hat{\bm{B}}_{\bm{a}}^{t}\bigr{)}^{-1}.$ (13)
@@
$\hat{x}_{i}^{t}(a_{i})=\sum_{k\in\mathcal{A}_{i}(a_{i})}\bigl{[}\hat{\bm{C}}_{\bm{a}}^{t}\bigr{]}_{ki}\epsilon_{k}+\sum_{l\not\in\mathcal{A}_{i}(a_{i})}\bigl{[}\hat{\bm{C}}_{\bm{a}}^{t}\bigr{]}_{li}\epsilon_{l},$ (14)
Since we do not assume any knowledge of the causal structure, all other nodes are considered as potential parents of node $i$. Based on (12) and (14), the residual can be expressed as
$r_{i}^{t}(a_{i})=-\epsilon_{i}+\\ \underbrace{\sum_{k\in\mathcal{A}_{i}(a_{i})}\bigl{(}\bigl{[}\hat{\bm{C}}_{\bm{a}}^{t}\bigr{]}_{ki}-\left[\bm{C}_{\bm{a}}\right]_{ki}\bigr{)}\epsilon_{k}}_{\textstyle\text{intrinsic error}}+\underbrace{\sum_{l\not\in\mathcal{A}_{i}(a_{i})}\bigl{[}\hat{\bm{C}}_{\bm{a}}^{t}\bigr{]}_{li}\epsilon_{l}}_{\textstyle\text{causal error}}.$ (15)
@@
$I(r_{i}^{t}(a_{i});x_{j})\leq\underbrace{I\Bigl{(}\sum_{l\not\in\mathcal{A}_{i}(a_{i})}\bigl{[}\hat{\bm{C}}_{\bm{a}}^{t}\bigr{]}_{li}\epsilon_{l}-\epsilon_{i};\sum_{l\not\in\mathcal{A}_{i}(a_{i})}\left[\bm{C}_{\bm{a}}\right]_{lj}\epsilon_{l}\Bigr{)}}_{\textstyle\text{causal error induced MI}}$
$+\underbrace{I\Bigl{(}\sum_{k\in\mathcal{A}_{i}(a_{i})}\bigl{(}\bigl{[}\hat{\bm{C}}_{\bm{a}}^{t}\bigr{]}_{ki}-\left[\bm{C}_{\bm{a}}\right]_{ki}\bigr{)}\epsilon_{k};\sum_{k\in\mathcal{A}_{i}(a_{i})}\left[\bm{C}_{\bm{a}}\right]_{kj}\epsilon_{k}\Bigr{)}}_{\textstyle\text{intrinsic error induced MI}}$ (17)
@@
The proof relies on the data processing inequality [31] and the mutual independence of the exogenous variables. We define the first and second parts on the right-hand side of (2) as the mutual information induced by the causal and intrinsic errors, respectively. Due to the existence of the intrinsic error, we may not be able to detect the causal error and thus misjudge causal directions. To illustrate this point, consider the causal error induced mutual information. Denote the linear combination of exogenous variables from non-ancestors by
@@
which is Gaussian, with mean $\mu_{\theta x}$ and variance $\sigma^{2}_{\theta x}$. As we have a causal error, there is an assumed causal edge that does not exist in the true graph. This erroneously causal edge, $j\to i$ (from non-ancestors) is used in the estimation, and as a result, induces a causal error in the residual, weighted by the estimated parameter. To focus on the effect caused by the erroneously labeled causal edge $j\rightarrow i$, we denote the causal error in the residual induced by other nodes by
$\theta(r_{i}^{t}(a_{i}),\backslash j)\doteq\sum_{l\not\in\mathcal{A}_{i}(a_{i})}\bigl{[}\hat{\bm{C}}_{\bm{a}}^{t}\bigr{]}_{li}\epsilon_{l}-\epsilon_{i}-\bigl{[}\hat{\bm{B}}_{\bm{a}}^{t}\bigr{]}_{ji}\theta(x_{j},i),$ (19)
which is Gaussian as well, with mean $\mu_{\theta r}$ and variance $\sigma^{2}_{\theta r}$. In this way, we can rewrite and bound the causal error induced mutual information as
$I\Bigl{(}\sum_{l\not\in\mathcal{A}_{i}(a_{i})}\bigl{[}\hat{\bm{C}}_{\bm{a}}^{t}\bigr{]}_{li}\epsilon_{l}-\epsilon_{i};\sum_{l\not\in\mathcal{A}_{i}(a_{i})}\left[\bm{C}_{\bm{a}}\right]_{lj}\epsilon_{l}\Bigr{)}$
$\quad=I\Bigl{(}\bigl{[}\hat{\bm{B}}_{\bm{a}}^{t}\bigr{]}_{ji}\theta(x_{j},i)+\theta(r_{i}^{t}(a_{i}),\backslash j);\theta(x_{j},i)\Bigr{)}$ (20)
$\quad=h\Bigl{(}\bigl{[}\hat{\bm{B}}_{\bm{a}}^{t}\bigr{]}_{ji}\theta(x_{j},i)+\theta(r_{i}^{t}(a_{i}),\backslash j)\Bigr{)}+h\Bigl{(}\theta(x_{j},i)\Bigr{)}$
@@
$\quad\leq\log\Bigl{(}\bigl{(}|\bigl{[}\hat{\bm{B}}_{\bm{a}}^{t}\bigr{]}_{ji}|\sigma_{\theta x}+\sigma_{\theta r}\bigr{)}\sqrt{2\pi e}\Bigr{)}+\log\Bigl{(}\sigma_{\theta x}\sqrt{2\pi e}\Bigr{)}$
@@
$\quad=\log\Bigl{(}1+\frac{\sigma_{\theta x}}{\sigma_{\theta r}}|\bigl{[}\hat{\bm{B}}_{\bm{a}}^{t}\bigr{]}_{ji}|\Bigr{)}-\frac{1}{2}\log(1-\rho_{\theta}^{2}),$ (23)
@@
$\rho_{\theta}=\frac{\mathbb{E}[(\theta(x_{j},i)-\mu_{\theta x})\cdot(\theta(r_{i}^{t}(a_{i}),\backslash j)-\mu_{\theta r})]}{\sigma_{\theta x}\cdot\sigma_{\theta r}},$ (24)
and $h(\cdot)$ denotes the differential entropy. Note that the inequality (22) comes from the independence bound on entropy [31] and the Gaussianity of key random variables.
@@
$I_{\mathrm{w}}(r_{i}^{t}(a_{i});x_{j})\doteq I(r_{i}^{t}(a_{i});x_{j})-\log|\bigl{[}\hat{\bm{B}}_{\bm{a}}^{t}\bigr{]}_{ji}|.$ (25)
@@
$\exists(i,j):\quad(i,j)\in\mathcal{B},\quad(i,j)\notin\hat{\mathcal{B}},$ (26)
where $\hat{\mathcal{B}}$ denotes the estimated edge set. In the following subsections, we analyze the impact of FN and FP errors and propose an adaptive edge-rejecting strategy to minimize the FN error rate.
@@ III-C Analysis of Graph Errors
$\mathbb{E}\bigl{[}\bigl{[}\hat{\bm{B}}_{\bm{a}}^{t}\bigr{]}_{i,\hat{\mathcal{P}}_{i}(a_{i})}\big{|}\bm{X}^{1:t}_{\hat{\mathcal{P}}_{i}(a_{i})}(a_{i})\bigr{]}=\left[\bm{B}_{\bm{a}}\right]_{i,\hat{\mathcal{P}}_{i}(a_{i})}.$ (27)
@@
$\mathbb{E}\bigl{[}\bigl{[}\hat{\bm{B}}_{\bm{a}}^{t}\bigr{]}_{i,\hat{\mathcal{P}}_{i}(a_{i})}\big{|}\bm{X}^{1:t}_{\hat{\mathcal{P}}_{i}(a_{i})}(a_{i})\bigr{]}=\left[\bm{B}_{\bm{a}}\right]_{i,\hat{\mathcal{P}}_{i}(a_{i})}+$
$\hskip 16.0pt\bigl{[}\bm{X}^{1:t}_{\hat{\mathcal{P}}_{i}(a_{i})}(a_{i})\bm{X}^{1:t}_{\hat{\mathcal{P}}_{i}(a_{i})}(a_{i})^{T}\bigr{]}^{-1}\mathbb{E}\bigl{[}\bm{X}^{1:t}_{\hat{\mathcal{P}}_{i}(a_{i})}(a_{i})\cdot$
$\hskip 32.0pt\bm{X}^{1:t}_{\mathcal{P}_{i}\backslash\hat{\mathcal{P}}_{i}(a_{i})}(a_{i})^{T}\big{|}\bm{X}^{1:t}_{\hat{\mathcal{P}}_{i}(a_{i})}(a_{i})\bigr{]}\left[\bm{B}_{\bm{a}}\right]_{i,\mathcal{P}_{i}\backslash\hat{\mathcal{P}}_{i}(a_{i})}.$ (28)
@@
Observe that with only FP errors, the estimated edge weights are asymptotically unbiased. Although the estimated weights for any non-existent parents in $\hat{\mathcal{P}}_{i}(a_{i})$ is non-zero for a finite number of samples, the bias converges to zero as more samples are collected. In contrast, with the existence of FN errors, the estimated weights for both correctly identified and erroneously rejected parents are biased, resulting in inaccurate weights even with an infinite amount of data. Note that the bias is a function of the correlation between identified parents and rejected parents.
@@ III-D Adaptive Edge Rejection
1:  Initialize the edge set to include all possible directed edges: ^=$\hat{\mathcal{B}}=\{(i,j):\forall i,j\in\mathcal{V}\}$.
2:  while $(\mathcal{V},\hat{\mathcal{B}})$ is not a DAG do
3:     Compute edge weights $[\hat{\bm{B}}]_{ij}$ and residuals, $\forall(i,j)\in\hat{\mathcal{B}}$, by MMSE estimation, with observed values $\bm{X}^{1:t}$.
4:     Compute edge-weighted MI $I_{\mathrm{w}}(r_{i}^{t}(0);x_{j})$, $\forall(i,j)\in\hat{\mathcal{B}}$, with $\bm{X}^{1:t}$ and empirical MI.
5:     Find $(i,j)=\operatorname*{arg\,max}_{i,j}I_{\mathrm{w}}(r_{i}^{t}(0);x_{j})$, remove the edge $(i,j)$ from $\hat{\mathcal{B}}$.
@@
7:  Compute $[\hat{\bm{B}}]_{ij}$, $\forall(i,j)\in\hat{\mathcal{B}}$, by MMSE estimation, with $\bm{X}^{1:t}$. For $(i,j)\notin\hat{\mathcal{B}}$, set $[\hat{\bm{B}}]_{ij}=0$.
8:  return  Estimated weight matrix $\hat{\bm{B}}$.

### v2

@@ III-B Edge-weighted Mutual Information
As a general measure of dependence, mutual information (MI), denoted by $I(\cdot)$, is widely utilized for causal discovery (see e.g. [35, 36]). However, since the sub-graphs are treated separately, pure MI-based sub-graph learning with limited data suffers significantly from estimation error. In this subsection, we explain this issue in detail and propose the edge-weighted mutual information measure as a solution.
@@
For example, we have $x_{3}=2.76\epsilon_{1}+2.3\epsilon_{2}+\epsilon_{3}$ for node $3$ in Fig. 1. With the estimated edge-weight matrix denoted by $\widehat{\bm{B}}_{\bm{a}}^{t}$, the estimated flow-weight matrix is defined as
$\widehat{\bm{C}}_{\bm{a}}^{t}\doteq\bigl{(}\bm{I}-\widehat{\bm{B}}_{\bm{a}}^{t}\bigr{)}^{-1}.$ (13)
@@
$\hat{x}_{i}^{t}=\sum_{k\in\mathcal{A}_{i}(a_{i})}\bigl{[}\widehat{\bm{C}}_{\bm{a}}^{t}\bigr{]}_{ki}\epsilon_{k}+\sum_{l\not\in\mathcal{A}_{i}(a_{i})}\bigl{[}\widehat{\bm{C}}_{\bm{a}}^{t}\bigr{]}_{li}\epsilon_{l}+\nu_{i}.$ (14)
For example, the estimate of $x_{3}$ can be decomposed according to the ancestor set of node $3$, as presented in Fig. 1. Since we do not assume any knowledge of the causal structure, all other nodes are considered as potential parents of node $i$. Based on (12) and (14), the residual can be expressed as
$r_{i}^{t}(a_{i})=-\tilde{\epsilon}_{i}+\\ \underbrace{\sum_{k\in\mathcal{A}_{i}(a_{i})}\bigl{(}\bigl{[}\widehat{\bm{C}}_{\bm{a}}^{t}\bigr{]}_{ki}-\left[\bm{C}_{\bm{a}}\right]_{ki}\bigr{)}\epsilon_{k}}_{\textstyle\text{intrinsic error}}+\underbrace{\sum_{l\not\in\mathcal{A}_{i}(a_{i})}\bigl{[}\widehat{\bm{C}}_{\bm{a}}^{t}\bigr{]}_{li}\epsilon_{l}}_{\textstyle\text{causal error}}.$ (15)
@@
$I(r_{i}^{t}(a_{i});x_{j})\leq\underbrace{I\Bigl{(}\sum_{l\not\in\mathcal{A}_{i}(a_{i})}\bigl{[}\widehat{\bm{C}}_{\bm{a}}^{t}\bigr{]}_{li}\epsilon_{l}-\tilde{\epsilon}_{i};\sum_{l\not\in\mathcal{A}_{i}(a_{i})}\left[\bm{C}_{\bm{a}}\right]_{lj}\epsilon_{l}\Bigr{)}}_{\textstyle\text{causal errorinduced MI}}$
$+\underbrace{I\Bigl{(}\sum_{k\in\mathcal{A}_{i}(a_{i})}\bigl{(}\bigl{[}\widehat{\bm{C}}_{\bm{a}}^{t}\bigr{]}_{ki}-\left[\bm{C}_{\bm{a}}\right]_{ki}\bigr{)}\epsilon_{k};\sum_{k\in\mathcal{A}_{i}(a_{i})}\left[\bm{C}_{\bm{a}}\right]_{kj}\epsilon_{k}\Bigr{)}}_{\textstyle\text{intrinsic error induced MI}}$ (17)
@@
The proof relies on the data processing inequality [37] and the mutual independence of the exogenous variables. We define the first and second terms on the right-hand side of (2) as the mutual information induced by the causal and intrinsic errors, respectively. Due to the existence of the intrinsic error, we may not be able to detect the causal error and thus misjudge causal directions. To illustrate this point, consider the causal error induced mutual information. Denote the linear combination of exogenous variables from non-ancestors by
@@
which is Gaussian, with mean $\mu_{\theta x}$ and variance $\sigma^{2}_{\theta x}$. For example, since the ancestor set of node $3$ in Fig. 1 is $\{1,2\}$, $\theta(x_{5},3)$ is a linear combination of $\epsilon_{3}$, $\epsilon_{4}$ and $\epsilon_{5}$. As we have a causal error, there is an assumed causal edge that does not exist in the true graph. This erroneous causal edge, $j\to i$ (from non-ancestors) is used in the estimation, and as a result, induces a causal error in the residual, weighted by the estimated parameter. To focus on the effect caused by the erroneously labeled causal edge $j\rightarrow i$, we denote the causal error in the residual induced by other nodes by
$\theta(r_{i}^{t}(a_{i}),\backslash j)\doteq\sum_{l\not\in\mathcal{A}_{i}(a_{i})}\bigl{[}\widehat{\bm{C}}_{\bm{a}}^{t}\bigr{]}_{li}\epsilon_{l}-\tilde{\epsilon}_{i}-\bigl{[}\widehat{\bm{B}}_{\bm{a}}^{t}\bigr{]}_{ji}\theta(x_{j},i),$ (19)
which is Gaussian as well, with mean $\mu_{\theta r}$ and variance $\sigma^{2}_{\theta r}$, while \$\backslash j$ denotes the set of other non-ancestors, \$\backslash j\doteq\{l\in\{1,\dots,N\}|l\notin\mathcal{A}_{i}(a_{i}),l\neq j\}$. In this way, we can rewrite and bound the causal error induced mutual information as
$I\Bigl{(}\sum_{l\not\in\mathcal{A}_{i}(a_{i})}\bigl{[}\widehat{\bm{C}}_{\bm{a}}^{t}\bigr{]}_{li}\epsilon_{l}-\tilde{\epsilon}_{i};\sum_{l\not\in\mathcal{A}_{i}(a_{i})}\left[\bm{C}_{\bm{a}}\right]_{lj}\epsilon_{l}\Bigr{)}$
$\quad=I\Bigl{(}\bigl{[}\widehat{\bm{B}}_{\bm{a}}^{t}\bigr{]}_{ji}\theta(x_{j},i)+\theta(r_{i}^{t}(a_{i}),\backslash j);\theta(x_{j},i)\Bigr{)}$ (20)
$\quad=h\Bigl{(}\bigl{[}\widehat{\bm{B}}_{\bm{a}}^{t}\bigr{]}_{ji}\theta(x_{j},i)+\theta(r_{i}^{t}(a_{i}),\backslash j)\Bigr{)}+h\Bigl{(}\theta(x_{j},i)\Bigr{)}$
@@
$\quad\leq\log\Bigl{(}\bigl{(}|\bigl{[}\widehat{\bm{B}}_{\bm{a}}^{t}\bigr{]}_{ji}|\sigma_{\theta x}+\sigma_{\theta r}\bigr{)}\sqrt{2\pi e}\Bigr{)}+\log\Bigl{(}\sigma_{\theta x}\sqrt{2\pi e}\Bigr{)}$
@@
$\quad=\log\Bigl{(}1+\frac{\sigma_{\theta x}}{\sigma_{\theta r}}|\bigl{[}\widehat{\bm{B}}_{\bm{a}}^{t}\bigr{]}_{ji}|\Bigr{)}-\frac{1}{2}\log(1-\rho_{\theta}^{2}),$ (23)
@@
$\rho_{\theta}=\frac{\mathbb{E}[(\theta(x_{j},i)-\mu_{\theta x})(\theta(r_{i}^{t}(a_{i}),\backslash j)-\mu_{\theta r})]}{\sigma_{\theta x}\sigma_{\theta r}},$ (24)
and $h(\cdot)$ denotes the differential entropy. Note that the inequality (22) comes from the Gaussianity of key random variables and the fact that the variance is maximized when the variables are perfectly correlated ($\rho_{\theta}=1$).
@@
$I_{\mathrm{w}}(r_{i}^{t}(a_{i});x_{j})\doteq I(r_{i}^{t}(a_{i});x_{j})-\log|\bigl{[}\widehat{\bm{B}}_{\bm{a}}^{t}\bigr{]}_{ji}|.$ (25)
@@
$\exists(i,j):\quad(i,j)\in\mathcal{B},\quad(i,j)\notin\widehat{\mathcal{B}},$ (26)
where $\widehat{\mathcal{B}}$ denotes the estimated edge set. In the following subsections, we analyze the impact of FN and FP errors and propose an adaptive edge-rejecting strategy to minimize the FN error rate.
@@ III-C Analysis of Graph Errors
$\mathbb{E}\bigl{[}\bigl{[}\widehat{\bm{B}}_{\bm{a}}^{t}\bigr{]}_{i,\widehat{\mathcal{P}}_{i}(a_{i})}\big{|}\bm{X}^{1:t}_{\widehat{\mathcal{P}}_{i}(a_{i})}(a_{i})\bigr{]}=\left[\bm{B}_{\bm{a}}\right]_{i,\widehat{\mathcal{P}}_{i}(a_{i})}.$ (27)
@@
$\mathbb{E}\bigl{[}\bigl{[}\widehat{\bm{B}}_{\bm{a}}^{t}\bigr{]}_{i,\widehat{\mathcal{P}}_{i}(a_{i})}\big{|}\bm{X}^{1:t}_{\widehat{\mathcal{P}}_{i}(a_{i})}(a_{i})\bigr{]}=\left[\bm{B}_{\bm{a}}\right]_{i,\widehat{\mathcal{P}}_{i}(a_{i})}+$
$\hskip 16.0pt\Bigl{[}\bm{X}^{1:t}_{\widehat{\mathcal{P}}_{i}(a_{i})}(a_{i}){\bm{X}^{1:t}_{\widehat{\mathcal{P}}_{i}(a_{i})}(a_{i})}^{\top}\Bigr{]}^{-1}\mathbb{E}\Bigl{[}\bm{X}^{1:t}_{\widehat{\mathcal{P}}_{i}(a_{i})}(a_{i})\cdot$
$\hskip 32.0pt{\bm{X}^{1:t}_{\mathcal{P}_{i}\backslash\widehat{\mathcal{P}}_{i}(a_{i})}(a_{i})}^{\top}\Big{|}\bm{X}^{1:t}_{\widehat{\mathcal{P}}_{i}(a_{i})}(a_{i})\Bigr{]}\left[\bm{B}_{\bm{a}}\right]_{i,\mathcal{P}_{i}\backslash\widehat{\mathcal{P}}_{i}(a_{i})}.$ (28)
@@
Observe that with only FP errors, the estimated edge weights are asymptotically unbiased. Although the estimated weights for any non-existent parents in $\widehat{\mathcal{P}}_{i}(a_{i})$ is non-zero for a finite number of samples, the bias converges to zero as more samples are collected. In contrast, with the existence of FN errors, the estimated weights for both correctly identified and erroneously rejected parents are biased, resulting in inaccurate weights even with an infinite amount of data. Note that the bias is a function of the correlation between identified parents and rejected parents.
@@ III-D Adaptive Edge Rejection
1:  Initialize the edge set to include all possible directed edges: ^=$\widehat{\mathcal{B}}=\{(i,j):\forall i,j\in\mathcal{V}\}$.
2:  while $(\mathcal{V},\widehat{\mathcal{B}})$ is not a DAG do
3:     Compute edge weights $\widehat{\bm{B}}_{ij}$ and residuals, $\forall(i,j)\in\widehat{\mathcal{B}}$, by MMSE estimation, with observed values $\bm{X}^{1:t}$.
4:     Compute edge-weighted MI $I_{\mathrm{w}}(r_{i}^{t}(0);x_{j})$, $\forall(i,j)\in\widehat{\mathcal{B}}$, with $\bm{X}^{1:t}$ and empirical MI.
5:     Find $(i,j)=\operatorname*{arg\,max}_{i,j}I_{\mathrm{w}}(r_{i}^{t}(0);x_{j})$, remove the edge $(i,j)$ from $\widehat{\mathcal{B}}$.
@@
7:  Compute $\widehat{\bm{B}}_{ij}$, $\forall(i,j)\in\widehat{\mathcal{B}}$, by MMSE estimation, with $\bm{X}^{1:t}$. For $(i,j)\notin\widehat{\mathcal{B}}$, set $\widehat{\bm{B}}_{ij}=0$.
8:  return  Estimated weight matrix $\widehat{\bm{B}}$.
@@
Lastly, we provide the time and space complexity of the proposed algorithm. Starting with a set of all possible edges of size $N^{2}$, the algorithm rejects an edge in each step until the graph becomes a DAG, and thus the number of steps is less than $N^{2}$. Each step selects an edge with the largest edge-weighted MI among the remaining edges, with time complexity $\mathcal{O}(N^{2})$. After the rejection of an edge, estimation of the weights and residuals takes $\mathcal{O}(N^{2}(N+t))$ time while estimation of the MI takes $\mathcal{O}(Nt\log t)$ time [38]. Besides, checking whether the graph is a DAG by topological sorting takes $\mathcal{O}(N^{2})$ time [39]. Finally, if $N\leq t$, the time complexity of the proposed scheme is $\mathcal{O}(N^{4}t+N^{3}t\log t)$. The space complexity is $\mathcal{O}(Nt)$ for storing the collected samples and a matrix of edge-weighted MI.